Mark is 16 years old. In 4 years he will be twice Norma’s age. How old is Norma now?
Found 3 solutions by Edwin McCravy, ikleyn, Alan3354:
Answer by Edwin McCravy(20056)

Look at the question first to see what the unknown is:

How old is Norma now?
Morma's age = N

Mark is 16 years old.
Mark's age = 16

In 4 years he will be twice Norma’s age.
Mark's age in 4 years = 16+4 = 20
Norma's age in 4 years = N+4

Look at this sentence again:

In 4 years he will be twice Norma’s age.
20 = 2(N+4)

You finish.

Edwin

Answer by ikleyn(52784)

.
If Mark is 16 years now, he will be 20 years old in 4 years.


Since Mark will be 20 years old in 4 years, Norma will be 10 years old at that time.


Hence, Norma is 10-4 = 6 years old now.


Answer.  Norma is 6 years old now.

It is a 2-nd grade problem, which can be solved (and, I believe, MUST be solved) without using any equations.


===============

This my note goes specially to Alan:

    There is NO an ambiguity in the given formulation.


    As formulated, it ALWAYS means "the age then".


    If the opposite is not pointed directly and explicitly.


Answer by Alan3354(69443)

Mark is 16 years old. In 4 years he will be twice Norma’s age. How old is Norma now?
-------------
It's ambiguous.
Will he be 2x Norma's age now, or her age then?